Calculate arithmetic mean with the help of following data. Use step deviation method.

Wages (in Rs.)

20

25

30

40

42

50

No. of students

4

6

10

5

5

10



1
Expert's answer
2021-06-29T10:25:23-0400
"20\times4=80"

"25\times6=150"

"30\times10=300"

"40\times5=200"

"42\times5=210"

"50\times10=500"

"80+150+300+200+210+500=1440"

"4+6+10+5+5+10=40"

"\frac{1440}{40}=36"